Hello and thanks for your help. Find the slope and y intercept of the line represeted by the following equations. 2x-3y=6 Mcgraw Hill 6th edition. Ch. 7 Ex. 1 Problem # 12. Thanks so much!!
Answer by checkley71(8403)

2X-3Y=6 OR -3Y=-2X+6 OR Y=2/3X-2 THUS A SLOPE OF 2/3 & THE Y INTERCEPT OF -2